Find equations in terms of y=mx+b in the following:
through (5, -2) and (1,3)

Answer by checkley71(8403)

Y=mX+b
(m)=SLOPE=(Y2-Y1)/(X2-X1)=(3+2)/(1-5)=5/-4=-5/4 FOR THE SLOPE.
NOW REPLACE THE X & Y UNKNOWNS BY ONE SET OF VALUES & SOLVE FOR THE Y INTERCEPT
(b).
3=-5/4*1+b
3=-5/4+b
b=3+5/4
b=(3*4+5)/4
b=(12+5)/4
b=17/4 ANSWER. THEREFORE THE EQUATION IS:
Y=-5X/4+17/4